# Neural Conditional Probability for Inference

### v1

@@ Abstract
We introduce NCP (Neural Conditional Probability), a novel operator-theoretic approach for learning conditional distributions with a particular focus on inference tasks. NCP can be used to build conditional confidence regions and extract important statistics like conditional quantiles, mean, and covariance. It offers streamlined learning through a single unconditional training phase, facilitating efficient inference without the need for retraining even when conditioning changes. By tapping into the powerful approximation capabilities of neural networks, our method efficiently handles a wide variety of complex probability distributions, effectively dealing with nonlinear relationships between input and output variables. Theoretical guarantees ensure both optimization consistency and statistical accuracy of the NCP method. Our experiments show that our approach matches or beats leading methods using a simple Multi-Layer Perceptron (MLP) with two hidden layers and GELU activations. This demonstrates that a minimalistic architecture with a theoretically grounded loss function can achieve competitive results without sacrificing performance, even in the face of more complex architectures.
@@ 1 Introduction
This paper addresses the task of estimating the conditional distribution $\mathbb{P}[Y|X]$ of the random variables $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ based on a dataset $\mathcal{D}_{n}:=(x_{i},y_{i})_{i\in[n]}$, consisting of observations sampled from their joint distribution. Learning conditional distributions is a fundamental problem in machine learning, crucial for various purposes such as building prediction intervals, performing downstream analysis, visualizing data, and interpreting outcomes. This entails predicting the probability of an event given certain conditions or variables, which is a crucial task across various domains, ranging from finance (Markowitz,, 1958) to medicine (Ray et al.,, 2017), to climate modeling (Harrington,, 2017) and beyond. For instance, in finance, it is essential for risk assessment to estimate the probability of default given economic indicators. Similarly, in healthcare, predicting the likelihood of a disease, given patient symptoms, aids in diagnosis. In climate modeling, estimating the conditional probability of extreme weather events such as hurricanes or droughts, given specific climate indicators, helps in disaster preparedness and mitigation efforts.
@@
According to Gao and Hastie, (2022), there exist four main strategies to learn the conditional distribution. The first one relies on the Bayes formula for densities and proposes to apply non-parametric statistics to learn the joint and marginal densities separately. However, most of non-parametric techniques face a significant challenge known as the curse of dimensionality (Scott,, 1991; Nagler and Czado,, 2016). The second strategy, also known as Localization method, involves training a model unconditionally on reweighted samples, where weights are determined by their proximity to the desired conditioning point (Hall et al.,, 1999; Yu and Jones,, 1998). However, these methods demand retraining the model whenever the conditioning changes and are prone to the curse of dimensionality if the weighting strategy treats all covariates equally. The third strategy, known as Direct Learning of the conditional distribution involves finding the best linear approximation of the conditional density on a dictionary of base functions or a kernel space (Sugiyama et al.,, 2010; Li et al.,, 2007). The performance of these methods relies crucially on the selection of bases and kernels. Again for high-dimensional settings, approaches that assign equal importance to all covariates may be less effective. Finally, the fourth strategy, known as Conditional Training, is an approach in which models are trained to estimate a target variable conditional on certain covariates or conditions. This typically involves partitioning the covariate space $\mathcal{X}$ into sets, followed by unconditional training of models for each set of the partition (see Gao and Hastie,, 2022; Winkler et al.,, 2020; Lu and Huang,, 2020; Dhariwal and Nichol,, 2021, and references therein). However, this strategy requires a large dataset to provide enough samples for each conditioning and is expensive as it requires training separate models for each conditioning input set, even though they stem from the same underlying joint distribution.
@@
Contributions   The principal contribution of this work is a different conditional probability approach that does not fall into any of the four aforementioned strategies. Our method, called Neural Conditional Probability (NCP), aims to learn the conditional expectation operator ${\mathsf{E}_{Y|X}}$ defined as $[{\mathsf{E}_{Y|X}}f](x):=\mathbb{E}[f(Y)\,|\,X=x]$, where $f:\mathcal{Y}\to\mathbb{R}$ is any measurable function. To this end, we introduce a principled loss to solve this optimization problem leveraging on recent results in (Kostic et al.,, 2024) concerning the relationship between the conditional expectation operator and deepCCA (Andrew et al.,, 2013). Once trained, this operator can be used interchangeably to:
@@
    estimate the conditional probabilities $\mathbb{P}[Y\in B|X\in A]$ for arbitrary sets $B\subset\mathcal{Y}$ and $A\subset\mathcal{X}$ with theoretical non-asymptotic guarantees on the accuracy of the estimates, meaning we can also easily build conditional confidence regions with our NCP method.
Notably, our approach extracts statistics directly from the trained operator without retraining or resampling, and it is supported by both optimization consistency and statistical guarantees. In addition our experiments show that our approach matches or exceeds the performance of leading methods, even when using a basic Multi-Layer Perceptron (MLP) with two hidden layers and GELU activations. This demonstrates the effectiveness of a minimalistic architecture combined with a theoretically grounded loss function.
Paper organization   In Section 2 we review related work. Section 3 introduces the operator theoretic approach to model conditional expectation, while Section 4 discusses its training pipeline. In Section 5, we derive theoretical non-asymptotic guarantees on the proposed method’s estimation error. Finally, Section 6 presents numerical experiments with the proposed method.
@@ 2 Related works
Non-parametric estimators are valuable for density and conditional density estimation as they don’t rely on specific assumptions about the density being estimated. Kernel estimators, pioneered by Parzen, (1962) and Rosenblatt, (1956), are a widely used non-parametric density estimation method. Much effort has been dedicated to enhancing kernel estimation, focusing on aspects like bandwidth selection Goldenshluger and Lepski, (2011), non-linear aggregation Rigollet and Tsybakov, (2007), and computational efficiency Langrené and Warin, (2020), as well as extending it to conditional densities Bertin et al., (2014). A comprehensive review of kernel estimators and their variants is provided in Silverman, (2017). See also (Tsybakov,, 2009) for a statistical analysis of their performance. However, most of non-parametric techniques face a significant challenge known as the curse of dimensionality Scott, (1991); Nagler and Czado, (2016), meaning that the required sample size for accurate estimation grows exponentially with the dimensionality of the data Silverman, (2017). Additionally, the computational complexity also increases exponentially with dimensionality Langrené and Warin, (2020).
@@
Examples of localization methods include the work by Hall et al., (1999) for conditional CDF estimation using local logistic regression and locally adjusted Nadaraya-Watson estimation, as well as conditional quantiles estimation via local pinball loss minimization in Yu and Jones, (1998). Examples of direct learning of the conditional distribution include Sugiyama et al., (2010) via decomposition on a dictionary of base functions. Similarly, Li et al., (2007) explores quantile regression in reproducing Hilbert kernel spaces.
@@
Conditional training is a popular approach which was adopted in numerous works, as in the recent work by Gao and Hastie, (2022) where a parametric exponential model for the conditional density $p_{\theta}(y|x)$ is trained using the Lindsey method within each bin of a partition of the space $\mathcal{X}$. This strategy has also been implemented in several prominent classes of generative models, including Normalizing Flow (NF) and Diffusion Models (DM) Tabak and Vanden-Eijnden, (2010); Dinh et al., (2014); Rezende and Mohamed, 2015a ; Sohl-Dickstein et al., (2015). These models work by mapping a simple probability distribution into a more complex one. Conditional training approaches for NF and DM have been developed in many works including e.g. Winkler et al., (2020); Lu and Huang, (2020); Dhariwal and Nichol, (2021). In efforts to lower the computational burden of conditional diffusion models, an alternative approach used heuristic approximations applied directly to unconditional diffusion models on computer vision related tasks (see e.g. Song et al.,, 2023; Zhang et al.,, 2023). However, the effectiveness of these heuristics in accurately mimicking the true conditional distributions remains uncertain. Another crucial aspect of these classes of generative models is that while the probability distribution is modelled explicitly, the computation of any relevant statistic, say $\mathbb{E}[Y|X]$ is left as an implicit problem usually solved by sampling from $p_{\theta}(y|x)$ and then approximating $\mathbb{E}[Y|X]$ via simple Monte-Carlo integration. As expected, this approach quickly becomes problematic as the dimension of the output space $\mathcal{Y}$ becomes large.
@@
Conformal Prediction (CP) is a popular model-agnostic framework for uncertainty quantification Vovk et al., (1999). Conditional Conformal Prediction (CCP) was later developed to handle conditional dependencies between variables, allowing in principle for more accurate and reliable predictions. See Lei and Wasserman, (2014); Romano et al., (2019); Chernozhukov et al., (2021); Gibbs et al., (2023) and the references cited therein. However, (CP) and (CCP) are not without limitations. The construction of these guaranteed prediction regions need to be recomputed from scratch for each value of the confidence level parameter and of the conditioning for (CCP). In addition, the produced confidence regions tend to be conservative.
@@ 3 Operator Approach to Probability Modeling
Consider a pair of random variables $X$ and $Y$ taking values in probability spaces $(\mathcal{X},\Sigma_{\mathcal{X}},\mu)$ and $(\mathcal{Y},\Sigma_{\mathcal{Y}},\nu)$, respectively, where $\mathcal{X}$ and $\mathcal{Y}$ are state spaces, $\Sigma_{\mathcal{X}}$ and $\Sigma_{\mathcal{Y}}$ are sigma algebras, and $\mu$ and $\nu$ are probability measures. Let $\rho$ be the joint probability measure of $(X,Y)$ from the product space $\mathcal{X}\times\mathcal{Y}$. We assume that $\rho$ is absolutely continuous w.r.t. to the product measure of its marginals, that is $\rho\ll\mu\times\nu$, and denote the corresponding density by $p=d\rho/d(\mu\times\nu)$, so that $\rho(dx,dy)=p(x,y)\mu(dx)\nu(dy)$.
@@
$p(dy\,|\,x):=p(x,y)\nu(dy),\,\text{ that is }\;p(B\,|\,x):=\mathbb{P}[Y\in B\,|\,X=x],x\in\mathcal{X},\,B\in\Sigma_{\mathcal{Y}}.$ (1)
Our approach is based on the simple fact that for every $B\in\Sigma_{\mathcal{Y}}$, we have that $\mathbb{P}[Y\in B\,|\,X=x]=\mathbb{E}[\mathbb{1}_{B}(Y)\,|\,X=x]$, where $\mathbb{1}_{B}$ denotes the characteristic function of set $B$. More broadly we address the above problem by studying the conditional expectation operator ${\mathsf{E}_{Y|X}}\colon L^{2}_{\nu}(\mathcal{Y})\to L^{2}_{\mu}(\mathcal{X})$, which is defined, for every $f\in L^{2}_{\nu}(\mathcal{Y})$ and $x\in\mathcal{X}$, as
@@
where $L^{2}_{\mu}(\mathcal{X})$ and $L^{2}_{\mu}(\mathcal{X})$ denotes the Hilbert spaces of functions that are square integrable w.r.t. to $\mu$ and $\nu$, respectively. One readily verifies that $\lVert{\mathsf{E}_{Y|X}}\rVert=1$ and ${\mathsf{E}_{Y|X}}\mathbb{1}_{\mathcal{Y}}=\mathbb{1}_{\mathcal{X}}$.
A prominent feature of the above operator is that its rank can reveal the independence of the random variables. That is, $X$ and $Y$ are independent random variables if and only if ${\mathsf{E}_{Y|X}}$ is rank one operator, in which case we have that ${\mathsf{E}_{Y|X}}=\mathbb{1}_{\mathcal{X}}\otimes\mathbb{1}_{\mathcal{Y}}$. This motivates one to consider the deflated operator ${\mathsf{D}_{Y|X}}={\mathsf{E}_{Y|X}}-\mathbb{1}_{\mathcal{X}}\otimes\mathbb{1}_{\mathcal{Y}}\colon L^{2}_{\nu}(\mathcal{Y})\to L^{2}_{\mu}(\mathcal{X})$, for which we have that $\lVert{\mathsf{D}_{Y|X}}\rVert\leq 1$ and
$\mathbb{E}[f(Y)\,|\,X=x]=[{\mathsf{E}_{Y|X}}f](x)=[{\mathsf{D}_{Y|X}}f](x)+\mathbb{E}[f(Y)],\,f\in L^{2}_{\nu}(\mathcal{Y}).$
For dependent random variables, the deflated operator is nonzero. In many important situations, such as when the conditional probability distribution is a.e. absolutely continuous w.r.t. to the target measure, that is $p(\cdot\,|\,x)\ll\nu$ for $\mu$-a.e. $x\in\mathcal{X}$, the operator ${\mathsf{E}_{Y|X}}$ is compact, and, hence, we can write the SVD of ${\mathsf{E}_{Y|X}}$ and ${\mathsf{D}_{Y|X}}$ as
${\mathsf{E}_{Y|X}}=\sum_{i\in\mathbb{N}_{0}}\sigma^{\star}_{i}\,u^{\star}_{i}\otimes v^{\star}_{i},\quad\text{ and }\quad{\mathsf{D}_{Y|X}}=\sum_{i\in\mathbb{N}}\sigma^{\star}_{i}\,u^{\star}_{i}\otimes v^{\star}_{i},$ (2)
@@
$p(x,y)=\textstyle{\sum_{i\geq 0}}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y)=1+\textstyle{\sum_{i\geq 1}}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y).$
The mild assumption that ${\mathsf{E}_{Y|X}}$ is a compact operator allows one to approximate it arbitrarily well with a (large enough) finite rank (empirical) operator. Choosing the operator norm as the measure of approximation error one from the Erchart-Young-Mirsky theorem concludes that the best approximation is given by the truncated SVD, that is for every $r\in\mathbb{N}$,
${\mathsf{D}_{Y|X}}\approx[\![{\mathsf{D}_{Y|X}}]\!]_{r}:=\textstyle{\sum_{i\in[r]}}\sigma^{\star}_{i}\,u^{\star}_{i}\otimes v^{\star}_{i},\quad\text{ and }\quad[\![{\mathsf{D}_{Y|X}}]\!]_{r}\in\textstyle{\operatorname*{\text{\rm arg\,min}}_{\operatorname*{\text{\rm rank}}(A)\leq r}}\lVert{\mathsf{D}_{Y|X}}-A\rVert,$
where the minimum is given by $\sigma^{\star}_{r}$, and the minimizer is unique whenever $\sigma^{\star}_{r+1}<\sigma^{\star}_{r}$. This leads to the approximation of the joint density w.r.t. marginals $p(x,y)\approx 1+\textstyle{\sum_{i\in[r]}}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y)$, which implies the prediction
$\mathbb{E}[f(Y)\,|\,X=x]{\approx}\mathbb{E}[f(Y)]{+}([\![{\mathsf{D}_{Y|X}}]\!]_{r}f)(x){=}\mathbb{E}[f(Y)]+\textstyle{\sum_{i\in[r]}}\,\sigma^{\star}_{i}\,u^{\star}_{i}(x)\mathbb{E}[f(Y)\,v^{\star}_{i}(Y)],$ (3)
as well as the estimation of the conditional probability measures
$\mathbb{P}[Y\in B\,|\,X=x]$ $\approx\mathbb{P}[Y\in B]+\textstyle{\sum_{i\in[r]}}\,\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,\mathbb{E}[v^{\star}_{i}(Y)\,\mathbb{1}_{B}(Y)],\;\text{ and }$
$\mathbb{P}[Y\in B\,|\,X\in A]$ ${=}\frac{{\langle\mathbb{1}_{A},{\mathsf{E}_{Y|X}}\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}{\approx}\mathbb{P}[Y\in B]{+}{\sum_{i\in[r]}}\,\sigma^{\star}_{i}\,\frac{\mathbb{E}[u^{\star}_{i}(X)\mathbb{1}_{A}(X)]}{\mathbb{P}[X\in A]}\,\mathbb{E}[v^{\star}_{i}(Y)\,\mathbb{1}_{B}(Y)],$ (4)
where the latter approximation error is bounded in the following lemma.

### v2

# Neural Conditional Probability for Uncertainty Quantification

@@ Abstract
We introduce Neural Conditional Probability (NCP), an operator-theoretic approach to learning conditional distributions with a focus on statistical inference tasks. NCP can be used to build conditional confidence regions and extract key statistics such as conditional quantiles, mean, and covariance. It offers streamlined learning via a single unconditional training phase, allowing efficient inference without the need for retraining even when conditioning changes. By leveraging the approximation capabilities of neural networks, NCP efficiently handles a wide variety of complex probability distributions. We provide theoretical guarantees that ensure both optimization consistency and statistical accuracy. In experiments, we show that NCP with a 2-hidden-layer network matches or outperforms leading methods. This demonstrates that a a minimalistic architecture with a theoretically grounded loss can achieve competitive results, even in the face of more complex architectures.
@@ 1 Introduction
This paper studies the problem of estimating the conditional distribution associated with a pair of random variables, given a finite sample from their joint distribution. This problem is fundamental in machine learning, and instrumental for various purposes such as building prediction intervals, performing downstream analysis, visualizing data, and interpreting outcomes. This entails predicting the probability of an event given certain conditions or variables, which is a crucial task across various domains, ranging from finance (Markowitz,, 1958) to medicine (Ray et al.,, 2017), to climate modeling (Harrington,, 2017) and beyond. For instance, in finance, it is essential for risk assessment to estimate the probability of default given economic indicators. Similarly, in healthcare, predicting the likelihood of a disease, given patient symptoms, aids in diagnosis. In climate modeling, estimating the conditional probability of extreme weather events such as hurricanes or droughts, given specific climate indicators, helps in disaster preparedness and mitigation efforts.
@@
According to Gao and Hastie, (2022), there exist four main strategies to learn the conditional distribution. The first one relies on the Bayes formula for densities and proposes to apply non-parametric statistics to learn the joint and marginal densities separately. However, most of non-parametric techniques face a significant challenge known as the curse of dimensionality (Scott,, 1991; Nagler and Czado,, 2016). The second strategy, also known as Localization method, involves training a model unconditionally on reweighted samples, where weights are determined by their proximity to the desired conditioning point (Hall et al.,, 1999; Yu and Jones,, 1998). These methods require retraining the model whenever the conditioning changes and may also suffer from the curse of dimensionality if the weighting strategy treats all covariates equally. The third strategy, known as Direct Learning of the conditional distribution involves finding the best linear approximation of the conditional density on a dictionary of base functions or a kernel space (Sugiyama et al.,, 2010; Li et al.,, 2007). The performance of these methods relies crucially on the selection of bases and kernels. Again for high-dimensional settings, approaches that assign equal importance to all covariates may be less effective. Finally, the fourth strategy, known as Conditional Training, involves training models to estimate a target variable conditioned on certain covariates. This is typically based on partitioning the covariates space $\mathcal{X}$ into sets, followed by training models unconditionally within each partition (see Gao and Hastie,, 2022; Winkler et al.,, 2020; Lu and Huang,, 2020; Dhariwal and Nichol,, 2021, and references therein). However, this strategy requires a large dataset to provide enough samples for each conditioning and is expensive as it requires training separate models for each conditioning input set, even though they stem from the same underlying joint distribution.
@@
Contributions   The principal contribution of this work is a different conditional probability approach that does not fall into any of the four aforementioned strategies. Rather than learning the conditional density directly, our method, called Neural Conditional Probability (NCP), aims to learn the conditional expectation operator ${\mathsf{E}_{Y|X}}$ associated to the random variables $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ based on data from their joint distribution. The operator is defined, for every measurable function $f:\mathcal{Y}\to\mathbb{R}$, as
$[{\mathsf{E}_{Y|X}}f](x):=\mathbb{E}[f(Y)\,|\,X=x].$
NCP is based on a principled loss, leveraging the connection between conditional expectation operators and deepCCA (Andrew et al.,, 2013) established in (Kostic et al.,, 2024), and can be used interchangeably to:
@@
    estimate the conditional probabilities $\mathbb{P}[Y\in B\,|\,X\in A]$ for arbitrary sets $B\subset\mathcal{Y}$ and $A\subset\mathcal{X}$ with theoretical non-asymptotic guarantees on accuracy, allowing us to easily construct conditional confidence regions.
Notably, our approach extracts statistics directly from the trained operator without retraining or resampling, and it is supported by both optimization consistency and statistical guarantees. In addition our experiments show that our approach matches or exceeds the performance of leading methods, even when using a basic a 2-hidden-layer network. This demonstrates the effectiveness of a minimalistic architecture combined with a theoretically grounded loss function.
Paper organization   In Section 2 we review related work. Section 3 introduces the operator theoretic approach to model conditional expectation, while Section 4 discusses its training pipeline. In Section 5, we derive learning guarantees for NCP. Finally, Section 6 presents numerical experiments.
@@ 2 Related works
Non-parametric estimators are valuable for density and conditional density estimation as they don’t rely on specific assumptions about the density being estimated. Kernel estimators, pioneered by Parzen, (1962) and Rosenblatt, (1956), are a widely used non-parametric density estimation method. Much effort has been dedicated to enhancing kernel estimation, focusing on aspects like bandwidth selection (Goldenshluger and Lepski,, 2011), non-linear aggregation (Rigollet and Tsybakov,, 2007), and computational efficiency (Langrené and Warin,, 2020), as well as extending it to conditional densities (Bertin et al.,, 2014). A comprehensive review of kernel estimators and their variants is provided in (Silverman,, 2017). See also (Tsybakov,, 2009) for a statistical analysis of their performance. However, most of non-parametric techniques face a significant challenge known as the curse of dimensionality (Scott,, 1991; Nagler and Czado,, 2016), meaning that the required sample size for accurate estimation grows exponentially with the dimensionality of the data (Silverman,, 2017). Additionally, the computational complexity also increases exponentially with dimensionality (Langrené and Warin,, 2020).
@@
Examples of localization methods include the work by Hall et al., (1999) for conditional CDF estimation using local logistic regression and locally adjusted Nadaraya-Watson estimation, as well as conditional quantiles estimation via local pinball loss minimization in (Yu and Jones,, 1998). Examples of direct learning of the conditional distribution include (Sugiyama et al.,, 2010) via decomposition on a dictionary of base functions. Similarly, Li et al., (2007) explores quantile regression in reproducing Hilbert kernel spaces.
@@
Conditional training is a popular approach which was adopted in numerous works, as in the recent work by Gao and Hastie, (2022) where a parametric exponential model for the conditional density $p_{\theta}(y|x)$ is trained using the Lindsey method within each bin of a partition of the space $\mathcal{X}$. This strategy has also been implemented in several prominent classes of generative models, including Normalizing Flow (NF) and Diffusion Models (DM) (Tabak and Vanden-Eijnden,, 2010; Dinh et al.,, 2014; Rezende and Mohamed, 2015a, ; Sohl-Dickstein et al.,, 2015). These models work by mapping a simple probability distribution into a more complex one. Conditional training approaches for NF and DM have been developed in many works including (e.g. Winkler et al.,, 2020; Lu and Huang,, 2020; Dhariwal and Nichol,, 2021). In efforts to lower the computational burden of conditional diffusion models, an alternative approach used heuristic approximations applied directly to unconditional diffusion models on computer vision related tasks (see e.g. Song et al.,, 2023; Zhang et al.,, 2023). However, the effectiveness of these heuristics in accurately mimicking the true conditional distributions remains uncertain. Another crucial aspect of these classes of generative models is that while the probability distribution is modelled explicitly, the computation of any relevant statistic, say $\mathbb{E}[Y|X]$ is left as an implicit problem usually solved by sampling from $p_{\theta}(y|x)$ and then approximating $\mathbb{E}[Y|X]$ via simple Monte-Carlo integration. As expected, this approach quickly becomes problematic as the dimension of the output space $\mathcal{Y}$ becomes large.
@@
Conformal Prediction (CP) is a popular model-agnostic framework for uncertainty quantification (Vovk et al.,, 1999). Conditional Conformal Prediction (CCP) was later developed to handle conditional dependencies between variables, allowing in principle for more accurate and reliable predictions (see Lei and Wasserman,, 2014; Romano et al.,, 2019; Chernozhukov et al.,, 2021; Gibbs et al.,, 2023, and the references cited therein). However, (CP) and (CCP) are not without limitations. The construction of these guaranteed prediction regions need to be recomputed from scratch for each value of the confidence level parameter and of the conditioning for (CCP). In addition, the produced confidence regions tend to be conservative.
@@ 3 Operator approach to probability modeling
Consider a pair of random variables $X$ and $Y$ taking values in probability spaces $(\mathcal{X},\Sigma_{\mathcal{X}},\mu)$ and $(\mathcal{Y},\Sigma_{\mathcal{Y}},\nu)$, respectively, where $\mathcal{X}$ and $\mathcal{Y}$ are state spaces, $\Sigma_{\mathcal{X}}$ and $\Sigma_{\mathcal{Y}}$ are sigma algebras, and $\mu$ and $\nu$ are probability measures. Let $\rho$ be the joint probability measure of $(X,Y)$ from the product space $\mathcal{X}\times\mathcal{Y}$. We assume that $\rho$ is absolutely continuous w.r.t. to the product measure of its marginals, that is $\rho\ll\mu\times\nu$, and denote the corresponding density by $p=d\rho/d(\mu\times\nu)$, also called point-wise dependency in Tsai et al., (2020), so that $\rho(dx,dy)=p(x,y)\mu(dx)\nu(dy)$.
@@
$p(B\,|\,x):=\mathbb{P}[Y\in B\,|\,X=x],\quad x\in\mathcal{X},\,B\in\Sigma_{\mathcal{Y}}.$ (1)
Our approach is based on the simple fact that $p(B\,|\,x)=\mathbb{E}[\mathbb{1}_{B}(Y)\,|\,X=x]$, where $\mathbb{1}_{B}$ denotes the characteristic function of set $B$. More broadly we address the above problem by studying the conditional expectation operator ${\mathsf{E}_{Y|X}}\colon L^{2}_{\nu}(\mathcal{Y})\to L^{2}_{\mu}(\mathcal{X})$, which is defined, for every $f\in L^{2}_{\nu}(\mathcal{Y})$ and $x\in\mathcal{X}$, as
@@
where $L^{2}_{\mu}(\mathcal{X})$ and $L^{2}_{\nu}(\mathcal{Y})$ denotes the Hilbert spaces of functions that are square integrable w.r.t. to $\mu$ and $\nu$, respectively. One readily verifies that $\lVert{\mathsf{E}_{Y|X}}\rVert=1$ and ${\mathsf{E}_{Y|X}}\mathbb{1}_{\mathcal{Y}}=\mathbb{1}_{\mathcal{X}}$.
A prominent feature of the above operator is that its rank can reveal the independence of the random variables. That is, $X$ and $Y$ are independent random variables if and only if ${\mathsf{E}_{Y|X}}$ is a rank one operator, in which case we have that ${\mathsf{E}_{Y|X}}=\mathbb{1}_{\mathcal{X}}\otimes\mathbb{1}_{\mathcal{Y}}$. It is thus useful to consider the deflated operator ${\mathsf{D}_{Y|X}}={\mathsf{E}_{Y|X}}-\mathbb{1}_{\mathcal{X}}\otimes\mathbb{1}_{\mathcal{Y}}\colon L^{2}_{\nu}(\mathcal{Y})\to L^{2}_{\mu}(\mathcal{X})$, for which we have that
$[{\mathsf{E}_{Y|X}}f](x)=\mathbb{E}[f(Y)]+[{\mathsf{D}_{Y|X}}f](x),\quad f\in L^{2}_{\nu}(\mathcal{Y}).$ (2)
For dependent random variables, the deflated operator is nonzero. In many important situations, such as when the conditional probability distribution is a.e. absolutely continuous w.r.t. to the target measure, that is $p(\cdot\,|\,x)\ll\nu$ for $\mu$-a.e. $x\in\mathcal{X}$, the operator ${\mathsf{E}_{Y|X}}$ is compact, and, hence, we can write the SVD of ${\mathsf{E}_{Y|X}}$ and ${\mathsf{D}_{Y|X}}$ respectively as
${\mathsf{E}_{Y|X}}=\sum_{i=0}^{\infty}\sigma^{\star}_{i}\,u^{\star}_{i}\otimes v^{\star}_{i},\quad\text{ and }\quad{\mathsf{D}_{Y|X}}=\sum_{i=1}^{\infty}\sigma^{\star}_{i}\,u^{\star}_{i}\otimes v^{\star}_{i},$ (3)
@@
$p(x,y)=\textstyle{\sum_{i=0}^{\infty}}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y)=1+\textstyle{\sum_{i=1}^{\infty}}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y).$
The mild assumption that ${\mathsf{E}_{Y|X}}$ is a compact operator allows one to approximate it arbitrarily well with a (large enough) finite rank (empirical) operator. Choosing the operator norm as the measure of approximation error and appealing to the Eckart-Young-Mirsky Theorem (see Theorem 3 in Appendix B.1) one concludes that the best approximation is given by the truncated SVD, that is for every $d\in\mathbb{N}$,
${\mathsf{D}_{Y|X}}\approx[\![{\mathsf{D}_{Y|X}}]\!]_{d}:=\textstyle{\sum_{i=1}^{d}}\sigma^{\star}_{i}\,u^{\star}_{i}\otimes v^{\star}_{i},\quad\text{ and }\quad[\![{\mathsf{D}_{Y|X}}]\!]_{d}\in\operatorname*{\text{\rm arg\,min}}_{\operatorname*{\text{\rm rank}}(A)\leq d}\lVert{\mathsf{D}_{Y|X}}-A\rVert,$ (4)
where the minimum is given by $\sigma^{\star}_{d}$, and the minimizer is unique whenever $\sigma^{\star}_{d+1}<\sigma^{\star}_{d}$. This leads to the approximation of the joint density w.r.t. marginals $p(x,y)\approx 1+\sum_{i=1}^{d}\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,v^{\star}_{i}(y)$, so that
$\mathbb{E}[f(Y)\,|\,X=x]\approx\mathbb{E}[f(Y)]+\textstyle{\sum_{i=1}^{d}}\,\sigma^{\star}_{i}\,u^{\star}_{i}(x)\mathbb{E}[f(Y)\,v^{\star}_{i}(Y)],$ (5)
which in particular, choosing $f=\mathbb{1}_{B}$, gives
$\mathbb{P}[Y\in B\,|\,X=x]\approx\mathbb{P}[Y\in B]+\textstyle{\sum_{i=1}^{d}}\,\sigma^{\star}_{i}\,u^{\star}_{i}(x)\,\mathbb{E}[v^{\star}_{i}(Y)\,\mathbb{1}_{B}(Y)].$
Moreover, we have that
$\mathbb{P}[Y\in B\,|\,X\in A]$ $=$ $\frac{{\langle\mathbb{1}_{A},{\mathsf{E}_{Y|X}}\mathbb{1}_{B}\rangle}}{\mathbb{P}[X\in A]}\approx\mathbb{P}[Y\in B]+\sum_{i=1}^{d}\,\sigma^{\star}_{i}\,\frac{\mathbb{E}[u^{\star}_{i}(X)\mathbb{1}_{A}(X)]}{\mathbb{P}[X\in A]}\,\mathbb{E}[v^{\star}_{i}(Y)\,\mathbb{1}_{B}(Y)],$
for which the approximation error is bounded in the following lemma.